An car of mass m starts from rest and accelerates uniformly up a ramp. It reaches the top of the ramp in one second. The ramp has length L and makes an angle θ with the horizontal. Assume that all friction forces can be lumped together under a single force Ff associated with a single friction constant μk.

Find the acceleration of the car in terms of L.
Answer #1

Under uniform acceleration:

s=ut+1/2at2, where s is displacement,u is initial velocity, t is time interval, a is acceleration

Here,u= 0 m/s as car starts from rest, t= 1 sec (given in problem) , s= length of ramp=L.

So, L=0*1+1/2a(1)2 =>a = 2*L

So,acceleration in terms of L = 2*L

Note that information about mass of the object and coefficient of friction is not required in order to determine acceleration in terms of L.